# UniVCD: A New Method for Unsupervised Change Detection in the Open-Vocabulary Era

### v1

@@ Abstract
Index Terms:
Change Detection, Unsupervised Learning, Vision Foundation Models, Remote Sensing, Deep Learning.
@@ II-A Vision Foundation Models
At the same time, the development of multimodal learning has endowed many vision foundation models[clip, qwenvl, cogagent, florence] with cross-modal understanding capabilities. By associating visual and textual information, they can handle more complex tasks such as image captioning and visual question answering, further expanding the application scope of foundation models. Among purely visual foundation models, SAM2[sam2] is a powerful segmentation model with strong generalization and adaptability across diverse scenes. Among multimodal vision–language models, CLIP[clip] jointly trains image and text encoders and can effectively link visual content with semantic concepts.
@@ IV Experiment
UniVCD(orig.) 55.2 38.1 32.4 19.3 3.0 1.5 35.0 21.2 43.3 27.7 $\approx 0$ $\approx 0$
UniVCD(orig., postproc) 58.4 41.2 32.5 19.4 2.6 1.3 37.6 23.2 42.6 27.1 $\approx 0$ $\approx 0$
UniVCD-RemoteClip(orig.) 53.5 36.5 25.0 14.3 20.6 11.5 32.9 19.7 27.1 15.7 $\approx 0$ $\approx 0$
@@ IV-D Experimental Results
As shown in Table I, we compare UniVCD with a range of classical methods and recent unsupervised approaches on the BCD datasets. To assess performance under different input resolutions, we evaluate the model both on cropped $256\times 256$ patches (crop.) and on the original high-resolution images (orig.). Moreover, to evaluate the impact of the post-processing step described in Section 2, we report results both with (postproc) and without post-processing.
@@
Meanwhile, UniVCD is also applicable to change detection for small objects and for oblique photography, which are capabilities that many existing methods lack. Since current datasets rarely focus on these scenarios, we provide qualitative results in Appendix A.
@@ V Conclusion
Appendix A Other Features of UniVCD
@@
A-A Small-Object Change Detection
@@
A-B Change Detection on Oblique photography

### v2

@@ II-A Vision Foundation Models
At the same time, the development of multimodal learning has endowed many vision foundation models[clip, qwenvl, cogagent, florence] with cross-modal understanding capabilities. By associating visual and textual information, they can handle more complex tasks such as image captioning and visual question answering, further expanding the application scope of foundation models. Among purely visual foundation models, SAM2[sam2] is a powerful segmentation model with strong generalization and adaptability across diverse scenes. Among multimodal vision–language models, CLIP[clip] jointly trains image and text encoders and can effectively link visual content with semantic concepts. SAM3[sam3] further extends these capabilities by enabling exhaustive segmentation of all instances of an open-vocabulary concept specified by a short text phrase or exemplars.
@@ III-E Post-Processing for Change Detection and Segmentation
  3. 3.
    Optionally, we also experiment with directly using the prompt text as input to SAM3 to achieve even finer segmentation results.
@@ IV Experiment
SAM3-CD 36.9 22.6 22.1 12.4 9.6 5.0 30.9 18.2 42.1 26.7 21.1 11.8
UniVCD(orig.) 55.2 38.1 32.4 19.3 6.2 3.2 35.0 21.2 43.3 27.7 $\approx 0$ $\approx 0$
UniVCD(orig., postproc) 58.4 41.2 32.5 19.4 5.0 2.6 37.6 23.2 42.6 27.1 $\approx 0$ $\approx 0$
UniVCD(orig., SAM3 postproc) 60.4 43.2 31.9 18.9 15.2 8.2 39.8 24.9 43.7 28.0 $\approx 0$ $\approx 0$
@@ IV-D Experimental Results
As shown in Table I, we compare UniVCD with a range of classical methods and recent unsupervised approaches on the BCD datasets. To assess performance under different input resolutions, we evaluate the model both on cropped $256\times 256$ patches (crop.) and on the original high-resolution images (orig.). Moreover, to evaluate the impact of the post-processing step described in Section 2, we report results with and without post-processing (postproc), as well as results using the SAM3-based post-processing (sam3 postproc).
Meanwhile, we implement SAM3-CD, a baseline open-vocabulary change detection method directly based on SAM3. The specific implementation details can be found in Appendix A.
@@
Meanwhile, UniVCD is also applicable to change detection for small objects and for oblique photography, which are capabilities that many existing methods lack. Since current datasets rarely focus on these scenarios, we provide qualitative results in Appendix B.
@@ V Conclusion
Appendix A Implementation Details of SAM3-CD
The overall implementation of SAM3-CD is relatively concise and straightforward. Given a pair of temporal images $(I_{1},I_{2})$, the SAM3 model is first applied with a fixed confidence threshold $c$ to segment the target categories of interest, yielding the corresponding sets of masks $(M_{1},M_{2})$. Subsequently, the Intersection over Union (IoU) between masks from different temporal instances is computed, and mask pairs are matched and filtered based on their IoU discrepancies. In this manner, regions exhibiting significant changes are identified and extracted, leading to the final change detection results.
Appendix B Other Features of UniVCD
@@
B-A Small-Object Change Detection
@@
B-B Change Detection on Oblique photography